# GLiDRE: Generalist Lightweight model for Document-level Relation Extraction

### v1

@@ 3.3.1 Fully supervised setting
Table 1: Fully supervised results on the Re-DocRED dataset. We compare GLiDRE with strong DocRE models, LLM-based methods and GLiREL. F1 scores are reported on both the development and test sets. Results are taken from the original papers, except for DocRE models, which are reported following Gao et al. (2024).
@@
ATLOP (Zhou et al., 2021) $76.15_{\pm 0.23}$ $75.88_{\pm 0.23}$ $77.81_{\pm 0.71}$ $76.13_{\pm 0.28}$
KD-DocRE (Tan et al., 2022a) $77.88_{\pm 0.42}$ $77.12_{\pm 0.49}$ $78.28_{\pm 0.72}$ $77.60_{\pm 0.25}$
@@
LoRA FT LLaMA2-7B-Chat - - 53.0253.0253.02 52.7852.7852.78
LoRA FT LLaMA2-13B-Chat - - 52.4552.4552.45 52.1552.1552.15
LMRC LLaMA2-7B-Chat - - 72.9272.9272.92 72.3372.3372.33
LMRC LLaMA2-13B-Chat - - 74.6374.6374.63 74.0874.0874.08
LMRC LLaMA3-8B-Instruct - - 72.8072.8072.80 72.2172.2172.21
LMRC LLaMA3.1-8B-Instruct - - 72.9372.9372.93 72.3572.3572.35
Other GLiNER adaptation
GLiREL (Boylan et al., 2025) - - 54.1354.1354.13 53.2453.2453.24
GLiDRE (Ours) $77.76_{\pm 0.35}$ $76.70_{\pm 0.37}$ $77.83_{\pm 0.23}$ $76.80_{\pm 0.22}$
The results of our experiments presented in Table 1, demonstrate that GLiDRE achieves strong and competitive performance on the Re-DocRED benchmark. Our model attains a test F1 score of 77.8377.8377.83. This places GLiDRE in close competition with established DocRE models like ATLOP and KD-DocRE. While it does not surpass the current state-of-the-art methods like DREEAM and TTM-RE, it is important to note that these models incorporate additional mechanisms, such as evidence informations for DREEAM and the Turing Token Machine for TTM-RE that is designed to scale better than our approach with large datasets.
@@
3.3.2 Low-Resource Settings
In low-resource scenarios, we evaluate the efficiency or our model by comparing it against two powerful supervised baselines. We selected DREEAM, as it represents the state-of-the-art on the fully-supervised benchmark, and ATLOP, whose architecture and performances share similarities with our approach. This comparison is designed to determine the dataset size at which these conventional supervised methods can match the performance of our model and quantifying the advantages of our approach in data-scarce regimes.
Table 2: Few-shot F1 results for GLiDRE, ATLOP and DREEAM across various training set sizes ($N$). All models are trained on the same 5 subsets for each value of $N$. Best results per column are shown in bold.
Model N = 1 N = 5 N = 10 N = 50 N = 100 N = 500 N = 1000
GLiDRE $\mathbf{24.45}_{\pm 6.37}$ $\mathbf{33.71}_{\pm 6.11}$ $\mathbf{41.73}_{\pm 2.94}$ $\mathbf{54.54}_{\pm 1.39}$ $\mathbf{60.04}_{\pm 0.50}$ $69.26_{\pm 0.37}$ $\mathbf{72.09}_{\pm 0.31}$
The results presented in Table 2 clearly demonstrate the superiority of GLiDRE in data-scarce environments. In the extremely low-data regime (N $\leq$ 100), our model establishes a lead over both ATLOP and the state-of-the-art model, DREEAM. For instance, with only 10 training samples, GLiDRE achieves an F1 score of 41.7341.7341.73, surpassing DREEAM by a margin of over 14 F1 points. This significant advantage persists up to N=100 samples.
As the amount of training data increases to 500 and 1000 samples, the performance gap narrows, and the baselines become more competitive. This indicates that while conventional supervised methods eventually catch up, our model’s architecture provides a distinct and substantial advantage when labeled data is minimal.
@@ 3.3.3 Zero-Shot Setting
The zero-shot setting for Document-Level Relation Extraction remains a largely underexplored research area. Given the scarcity of existing baselines, we establish a benchmark by evaluating the performance of much larger open-weight Large Language Models. We selected good performing models from the Open LLM Leaderboard (Fourrier et al., 2024), specifically employing the instruction-tuned versions of Qwen 2.5 72B (Team, 2024), Mistral Large 123B [^3] https://huggingface.co/mistralai/Mistral-Large-Instruct-2411, Llama 3.3 70B [^4] https://huggingface.co/meta-llama/Llama-3.3-70B-Instruct and the model we used for synthetic pretraining generation, Mistral-Small 24B. Our prompting methodology is inspired by recent work in zero-shot Information Extraction (Yuan et al., 2023). We annotate entity mentions directly within the input text using special tags and to ensure structured and reliable outputs, we constrain the generation process to a strict (head, relation, tail) triplet format. This constrained decoding is implemented using regular expressions within the vLLM inference framework (Kwon et al., 2023), which facilitates parsing and mitigates the risk of hallucination. The detailed prompt template used for this task is provided in Figure 2. We use a fixed temperature of 0.
Our primary finding is that GLiDRE, a model with only a few hundred million parameters, achieves performance that is competitive with LLMs orders of magnitude larger. With a test F1 score of 17.3217.3217.32, GLiDRE outperforms Llama 3.3 70B and Mistral-Small 24B. This demonstrates that for specialized tasks like DocRE, a focused bi-encoder architecture can rival the zero-shot reasoning capabilities of massive general-purpose models. Furthermore, the results show a dramatic improvement when compared to the score of GPT-3.5 Turbo reported by Xue et al. (2024), showing the rapid recent progress of LLMs in zero-shot Information Extraction tasks.
Beyond performance, GLiDRE offers significant practical advantages. The LLM baselines entail substantial computational overhead, requiring vast memory footprints (e.g., over 300 GB for Mistral-Large and at least 4 A100-80GB) and significant energy consumption for inference. In contrast, GLiDRE operates with a memory footprint of under 10 GB. Moreover, LLMs often require complex engineering, including sophisticated prompting and constrained decoding, to produce structured, parsable outputs and mitigate hallucinations, while our model natively produces structured predictions.
The performance of GLiDRE shows some variability, which may come from a misalignment between their general-purpose pre-training corpora and the specific domain and relations of the Re-DocRED dataset. Adapting its pre-training by generating synthetic data with relation types aligned to the target dataset, as was done with GLiNER for Personally Identifiable Informations [^5] https://huggingface.co/urchade/gliner_multi_pii-v1, could improve results. However, this would likely reduce its ability to generalize to other DocRE datasets.
Table 3: Zero-shot results on the Re-DocRED dev and test sets. Results for GPT-3.5 Turbo are from Xue et al. (2024). All other LLMs are instruction-tuned versions.
@@
3.4 Analysis
3.4.1 LogSumExp vs. Mean Pooling
@@ 3.4 Analysis
However, our empirical results, shown in Table 4, indicate a different outcome within our architecture. For GLiDRE, conventional mean pooling outperforms LSE by nearly a full F1 point on the test set. We attribute this discrepancy to architectural differences; unlike ATLOP’s bilinear classifier, GLiDRE employs a bi-encoder framework that compares relation and type embeddings. This structural divergence suggests that optimizations are not always directly transferable between models. Given that the original ATLOP paper reported only minor gains from LSE, our findings confirm that mean pooling is a more effective and robust choice for our model.
@@
3.4.2 Effect of Pre-training and Localized Context Pooling
We conduct an ablation study to isolate the individual contributions of our synthetic pre-training stage and the Localized Context Pooling (noted LOP in the ATLOP model) mechanism. Two variants of our model are evaluated: one finetuned without pre-training and another without the LOP module.
The results in Table 5 confirm that both components positively contribute to the final performance of GLiDRE. Removing the pre-training stage leads to the most significant performance decrease, with a drop of nearly 0.70.70.7 F1 points, underscoring the effectiveness of our synthetic data generation for model initialization. Disabling LOP results in a drop of 0.20.20.2 F1 points, which validates its role in refining relation representations. These findings justify the inclusion of both techniques in our final model architecture.
Table 5: Ablation study on the Test set. We report F1 scores after removing the pre-training stage and the Localized Context Pooling (LOP) module.
@@ 3.4.2 Effect of Pre-training and Localized Context Pooling
3.4.3 Adapting the Adaptive Threshold Loss
@@
Adapting this to GLiDRE is non-trivial due to our model’s bi-encoder architecture, which lacks a fixed classifier head. To replicate the mechanism, we introduce a small multi-layer perceptron (MLP) to predict the threshold logit. The MLP input is a concatenation of the candidate relation’s embedding and a global context vector, formed by averaging all possible relation type embeddings. The model is then trained using the original ATLOP loss function.
To save compute resources, this experiment is conducted on the model variant without pre-training. As shown in Table 6, this adaptation proved detrimental, degrading performance by approximately 1.41.41.4 F1 points. We hypothesize several reasons for this negative result: (1) our method of computing the threshold logit via a separate MLP is fundamentally different from ATLOP’s integrated classifier approach; (2) the ATLOP loss may be less effective at handling the severe class imbalance in DocRED compared to the Focal Loss used in our main model; (3) a dynamic threshold may be unnecessary for our model. We observed that the optimal global threshold for GLiDRE consistently converges near 0.5 and further tuning per-class thresholds on the development set did not improve test set performance.
Table 6: Comparison between standard training with Focal Loss and our adaptation of the ATLOP adaptive thresholding method. Experiments are conducted without the pre-training phase.
@@
4 Conclusion
We present GLiDRE, a novel lightweight bi-encoder model for document-level relation extraction that reconceptualizes Document-level Relation Extraction as a direct representation matching problem. We show, through extensive experiments on the Re-DocRED benchmark, that GLiDRE achieves state-of-the-art few-shot performance and matches or surpasses much larger LLMs in zero-shot settings, all while operating with a fraction of their computational footprint.
GLiDRE not only rivals fully supervised baselines under low-resource regimes but also delivers highly structured predictions without complex prompting or constrained decoding. Its efficiency makes it a practical choice for real-world IE applications.
Future work will explore synthetic relation generation to further close remaining performance gaps and investigate dynamic thresholding methods tailored to the bi-encoder setup to enhance robustness across diverse relation schemas.
Limitations
Despite its efficiency and strong few-shot performance, GLiDRE faces several inherent limitations. First, the number of candidate relation pairs grows quadratically with the number of entities in a document, which can dramatically increase memory consumption for texts with high entity density and even lead to out-of-memory errors. Second, although the bi-encoder design can in principle accommodate longer contexts, it remains bound by the maximum sequence length of the underlying document encoder (e.g. 512 tokens for DeBERTa). Documents that exceed this limit must be truncated or segmented into chunks, potentially disrupting long-distance dependencies and harming performance.
Third, GLiDRE’s independent scoring of each entity-relation pair overlooks inter-label interactions. As a result, it can struggle to distinguish semantically similar relation types in the absence of a joint classification mechanism. Finally, our evaluation follows standard DocRE protocols by assuming gold entities and coreference chains are provided. Since GLiDRE does not perform named entity recognition or coreference resolution, its effectiveness in a fully end-to-end pipeline would depend on the accuracy of upstream modules, which may propagate errors and degrade overall performance.
Acknowledgments
Appendix A Appendix
A.1 Pre-training Data Generation
The synthetic data for the pre-training phase is generated using the prompt template detailed in Figure 3.
Figure 3: The prompt template used for pre-training data generation with Mistral-Small.
A.2 Zero-Shot Relation Extraction
@@
However, as the number of unseen relations increases, GLiDRE’s performance degrades more rapidly than models specifically architected for sentence-level zero-shot relation classification, such as GLiREL and TMC-BERT. We attribute this to our model’s pre-training on multi-mention instances, which may make it less specialized for the sentence-level relation classification. Furthermore, its bi-encoder design also scores each relation on its own, making it harder to tell apart similar relation types.

### v2

@@ 4.3.2 Few-Shot Document-level Relation Extraction
Table 3: Fully supervised setting: results on the Re-DocRED dataset. We compare GLiDRE with strong DocRE models, LLM-based methods and GLiREL. F1 scores are reported on both the development and test sets. Results are taken from the original papers, except for DocRE models, which are reported following Gao et al. (2024).
@@
ATLOP (Zhou et al., 2021) $76.15_{\pm 0.23}$ $5.88_{\pm 0.23} 77$$7.81_{\pm 0.71}$ $76.13_{\pm 0.28}$
KD-DocRE (Tan et al., 2022a) $7.88_{\pm 0.42} 77$$7.12_{\pm 0.49}$ $8.28_{\pm 0.72} 77$$7.60_{\pm 0.25}$
@@
LMRC LLaMA2-13B-Chat - - $74.63$ $74.08$
GLiNER-inspired model
GLiREL (Boylan et al., 2025) - - $54.13$ $53.24$
GLiDRE (Ours) $77.76_{\pm 0.35}$ $6.70_{\pm 0.37} 77$$7.83_{\pm 0.23}$ $76.80_{\pm 0.22}$
4.3.3 Fully supervised setting
Additionally, we evaluate the scalability of our proposed model, GLiDRE, by comparing it against several strong baselines on the Re-DocRED dataset under a fully supervised protocol. Specifically, in addition to DREEAM and ATLOP, we also include the following baselines:
KD-DocRE (Tan et al., 2022a) leverages an axial attention module to model entity-pairs interdependency across sentences, applies an adaptive focal loss to mitigate class imbalance and uses knowledge distillation to incorporate distantly supervised data.
TTM-RE (Gao et al., 2024) introduces a Token Turing Machine memory module that augments document representations with external memory tokens, paired with a noise‐robust loss tailored for positive–unlabeled setting.
Additionally, we benchmark against GLiREL (Boylan et al., 2025), a sentence-level relation extraction adaptation of GLiNER architecture. It is important to note that GLiREL was originally conceived for sentence-level relation extraction and operates with a shorter context length. For document-level predictions, it relies on aggregating relations based on gold-standard coreference information.
We also report recent methods that leverage LLMs, comparing with the framework introduced by Li et al. (2024b), which involves fine-tuning Llama models (Touvron et al., 2023) using Low-Rank Adaptation (Hu et al., 2022) and enhancing the model performance by employing a classifier to exhibit potential relations and guide the finetuned LLM. We only report their best result on this dataset.
The results presented in Table 3, demonstrate that GLiDRE achieves competitive performance on the Re-DocRED benchmark. Notably, this performance is attained despite the model not being specifically designed for large-scale datasets, highlighting its versatility and suitability across diverse scenarios. Our model attains a test F1 score of $77.83$. This places GLiDRE in close competition with established DocRE models like ATLOP and KD-DocRE. While it does not surpass the current state-of-the-art methods like DREEAM and TTM-RE, it is important to note that these models incorporate additional mechanisms, such as evidence information for DREEAM and the Turing Token Machine for TTM-RE that is designed to scale better than our approach with large datasets.
@@ 4.3.4 Zero-Shot Setting
Table 4: Zero-shot setting: results on the Re-DocRED dev and test sets. Results for GPT-3.5 Turbo are from Xue et al. (2024). All other LLMs are instruction-tuned versions.
@@
The zero-shot setting for Document-Level Relation Extraction remains a largely underexplored research area. Given the scarcity of existing baselines, we establish a benchmark by evaluating the performance of much larger open-weight Large Language Models. We compare against recent high-performing models, specifically employing the instruction-tuned versions of Qwen 2.5 72B (Team, 2024), Mistral Large 123B [^3] https://huggingface.co/mistralai/Mistral-Large-Instruct-2411, Llama 3.3 70B [^4] https://huggingface.co/meta-llama/Llama-3.3-70B-Instruct and the model we used for pretraining data generation, Mistral-Small 24B. Our prompting methodology is inspired by recent work in zero-shot Information Extraction (Yuan et al., 2023). We annotate entity mentions directly within the input text using special tags and, in order to ensure structured and reliable outputs, we constrain the generation process to a strict (head, relation, tail) triplet format. This constrained decoding is implemented using regular expressions within the vLLM inference framework (Kwon et al., 2023), which facilitates parsing and mitigates the risk of hallucination. The detailed prompt template used for this task is provided in Appendix A.2. We use a fixed temperature of 0.
Our primary finding is that GLiDRE, a model with only a few hundred million parameters, achieves performance that is competitive with LLMs orders of magnitude larger. With a test F1 score of $17.32$, GLiDRE outperforms Llama 3.3 70B and Mistral-Small 24B (which was used to generate the pretraining data) demonstrating the effectiveness to train on the pretraining corpus, as opposed to directly leveraging Mistral-Small 24B to annotate DocRED. This demonstrates that for specialized tasks like DocRE, a focused bi-encoder architecture can rival the zero-shot reasoning capabilities of massive general-purpose models. Furthermore, the results show a dramatic improvement when compared to the score of GPT-3.5 Turbo reported by Xue et al. (2024), showing the rapid recent progress of LLMs in zero-shot Information Extraction tasks.
Beyond performance, GLiDRE offers significant practical advantages. The inference process on the Re-DocRED test set (500 documents) with GLiDRE (0.8B parameters) requires less than 10GB of VRAM and completes in approximately 100 seconds on a single NVIDIA A100-80GB GPU. In contrast, inference with Mistral-Large (123B parameters) necessitates a distributed setup of at least 4x A100-80GB GPUs to accommodate its memory footprint of over 300GB. This process takes around 600 seconds and a total cost of approximately 2400 GPU-seconds. This breakdown demonstrates that GLiDRE is not only smaller but over 20 times more cost-effective in terms of total compute required for inference. Moreover, using LLMs for specific tasks often require complex engineering, including sophisticated prompting and constrained decoding, to produce structured, parsable outputs and mitigate hallucinations, while our model natively produces structured predictions.
The performance of GLiDRE shows some variability, which may come from a misalignment between their general-purpose pretraining corpora and the specific domain and relations of the Re-DocRED dataset. Adapting its pretraining by generating synthetic data with relation types aligned to the target dataset, as was done with GLiNER for Personally Identifiable Informations [^5] https://huggingface.co/urchade/gliner_multi_pii-v1, could improve results. However, this would likely reduce its ability to generalize to other DocRE datasets.
5 Conclusion
We present GLiDRE, a novel lightweight bi‑encoder model for document‑level relation extraction that reconceptualizes Document-level Relation Extraction as a direct representation matching problem. We show, through extensive experiments on the Re‑DocRED, FREDo and Re-FREDo benchmarks, that GLiDRE achieves state‑of‑the‑art few‑shot performance and matches or surpasses much larger LLMs in zero‑shot settings, all while operating with a fraction of their computational footprint.
GLiDRE not only rivals fully supervised baselines under low‑resource regimes but also delivers highly structured predictions without complex prompting or constrained decoding. Its efficiency makes it a practical choice for real‑world IE applications.
Future work will explore synthetic relation generation to further close remaining performance gaps and investigate dynamic thresholding methods tailored to the bi‑encoder setup to enhance robustness across diverse relation schemas.
Limitations
Despite its efficiency and strong few‑shot performance, GLiDRE faces several inherent limitations.
The number of candidate relation pairs grows quadratically with the number of entities in a document, which can dramatically increase memory consumption for texts with high entity density and even lead to out‑of‑memory errors.
Although the bi‑encoder design can in principle accommodate longer contexts, it remains bound by the maximum sequence length of the underlying document encoder (e.g. 512 tokens for DeBERTa). Other approaches we compare against, such as DREEAM, ATLOP, KD-DocRE, and TTM-RE, have the same limitation. Documents that exceed this limit must be truncated or segmented into chunks, potentially disrupting long‑distance dependencies and harming performance.
The independent scoring of each entity‐relation pair in GLiDRE overlooks inter‑label interactions. As a result, the model can struggle to distinguish semantically similar relation types in the absence of a joint classification mechanism.
Our evaluation follows standard DocRE protocols by assuming gold entities and coreference chains are provided. Since GLiDRE does not perform named entity recognition or coreference resolution, its effectiveness in a fully end‑to‑end pipeline would depend on the accuracy of upstream modules, which may propagate errors and degrade overall performance. This limitation is shared by other existing baselines and is particularly pronounced for DREEAM, which additionally relies on evidence annotations.
The pretraining of GLiDRE relies on the synthetic annotation of texts with a LLM, which is a source of risk for the propagation of the inherent biases of the LLM. This risk is mitigated by the use of clear annotation guidelines and could be further limited by combining multiple prompts and models and having human reviewers audit a subset of the annotations for bias and correctness.
Acknowledgments
Appendix A Appendix
A.1 Pretraining Data Generation
The synthetic data for the pretraining phase is generated using the prompt template detailed in Figure 2. An example of generated data is provided in Figure 3
Figure 2: The prompt template used for pretraining data generation with Mistral-Small.
Figure 3: Example of generated data.
A.2 Zero-shot DocRE inference with LLMs
@@
A.3 Analysis
A.3.1 LogSumExp vs. Mean Pooling
@@ A.3 Analysis
However, our empirical results, shown in Table 5, indicate a different outcome within our architecture. For GLiDRE, conventional mean pooling outperforms LSE by nearly a full F1 point on the test set. We attribute this discrepancy to architectural differences; unlike the bilinear classifier of ATLOP, GLiDRE employs a bi-encoder framework that compares relation and type embeddings. This structural divergence suggests that optimizations are not always directly transferable between models. Given that the original ATLOP paper reported only minor gains from LSE, our findings confirm that mean pooling is a more effective and robust choice for our model.
@@
A.3.2 Effect of pretraining and Localized Context Pooling
We conduct an ablation study to isolate the individual contributions of our synthetic pretraining stage and the Localized Context Pooling (noted LOP in the ATLOP model) mechanism. Two variants of our model are evaluated: one finetuned without pretraining and another without the LOP module.
The results in Table 6 confirm that both components positively contribute to the final performance of GLiDRE. Removing the pretraining stage leads to the most significant performance decrease, with a drop of nearly $0.7$ F1 points, underscoring the effectiveness of our synthetic data generation for model initialization. Disabling LOP results in a drop of $0.2$ F1 points, which validates its role in refining relation representations. These findings justify the inclusion of both techniques in our final model architecture.
Table 6: Ablation study on the Test set. We report F1 scores after removing the pretraining stage and the Localized Context Pooling (LOP) module.
@@ A.3.2 Effect of pretraining and Localized Context Pooling
A.3.3 Adapting the Adaptive Threshold Loss
@@
Adapting this to GLiDRE is non-trivial due to our bi-encoder architecture, which lacks a fixed classifier head. To replicate the mechanism, we introduce a small multi-layer perceptron (MLP) to predict the threshold logit. The MLP input is a concatenation of the candidate relation embedding and a global context vector, formed by averaging all possible relation type embeddings. The model is then trained using the original ATLOP loss function.
To save compute resources, this experiment is conducted on the model variant without pretraining. As shown in Table 7, this adaptation proved detrimental, degrading performance by approximately $1.4$ F1 points. We hypothesize several reasons for this negative result: (1) our method of computing the threshold logit via a separate MLP is fundamentally different from ATLOP’s integrated classifier approach; (2) the ATLOP loss may be less effective at handling the severe class imbalance in DocRED compared to the Focal Loss used in our main model; (3) a dynamic threshold may be unnecessary for our model. We observed that the optimal global threshold for GLiDRE consistently converges near 0.5 and further tuning per-class thresholds on the development set did not improve test set performance.
Table 7: Comparison between standard training with Focal Loss and our adaptation of the ATLOP adaptive thresholding method. Experiments are conducted without the pretraining phase.
@@
A.4 Zero-Shot Sentence-level Relation Extraction
@@
However, as the number of unseen relations increases, GLiDRE performance degrades more rapidly than models specifically architected for sentence-level zero-shot relation classification, such as GLiREL and TMC-BERT. We attribute this to our model pretraining on multi-mention instances, which may make it less specialized for the sentence-level relation classification. Furthermore, its bi‑encoder design also scores each relation on its own, making it harder to tell apart similar relation types.